# Towards Tight Convex Relaxations for Contact-Rich Manipulation

### v1

@@ VI-B The limit surface
and $\mu_{S}\in\mathbb{R}$ is the friction coefficient between the slider and table, $m_{S}\in\mathbb{R}$ is the slider mass, $g\in\mathbb{R}$ is the gravitional acceleration, $r\in\mathbb{R}$ is a characteristic distance, typically chosen as the max distance between a contact point and origin of frame $S$, and $c\in[0,1]$ is an integration constant that depends on the slider geometry, as detailed in [61] and [63]. From the principle of maximum dissipation, it can be shown that when the slider is sliding, the applied spatial contact force $F$ must lie on the limit surface, and the instantaneous spatial velocity $V$ must be perpendicular to the limit surface: $V\propto\nabla H(F)=DF$ , where ($\nabla$($\cdot$)∇ ( ⋅ ) denotes the gradient operator and $F$ and $V$ are measured in the same frame [47]. We note that this can be approximated as the linear constraint $V=DF$ and appropriately scaling $F$ as a post-processing step
@@ VI-C Slider-pusher dynamics
$R=\begin{bmatrix}\prescript{W}{}{R}^{S}&0\\ 0&1\\ \end{bmatrix},\quad\prescript{W}{}{R}^{S}\in SO(2)$
@@ VII-A Modes of the slider-pusher system
$r\coloneqq(c_{\theta},s_{\theta}),\quad\|r\|_{2}^{2}=1\iff\begin{bmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\\ \end{bmatrix}\in SO(2)$ (26)
@@ VII-B Motion planning within a mode
We discretize the dynamics in (16) using forward Euler discretization. We represent a trajectory segment within each mode for the slider-pusher system by discrete points for the state and input: $x_{0},x_{1},\ldots,x_{N}$ and $u_{0},u_{1},\ldots,u_{N-1}$. The dynamics are then $x_{k+1}=x_{k}+hg(x_{k},u_{k})$ for $k=0,\ldots,N-1$ with the chosen time step $h\in\mathbb{R}_{>0}$.
For the contact modes, we minimize the arc length, trajectory energy, contact forces, and time in contact through the following cost:
$\sum_{k=0}^{N-1}L(x_{k},x_{k+1})+E(x_{k},x_{k+1})+k_{f}\|f_{k}\|_{2}^{2}+k_{T}T$ (27)
where
$L(x_{k},x_{k+1})$ $=k_{p}(1/N_{v})\sum_{i=1}^{N_{v}}\|p_{k+1}^{v_{i}}-p_{k}^{v_{i}}\|_{2},$
$E(x_{k},x_{k+1})$ $=k_{v}(1/N_{v})\sum_{i=1}^{N_{v}}\|p_{k+1}^{v_{i}}-p_{k}^{v_{i}}\|_{2}^{2}$
@@
models the arc length on $\mathrm{SE}(2)$ and energy of the slider, respectively, and $N_{v}$ is the number of slider vertices. $L(x_{k},x_{k+1})$ corresponds to the mean distance travelled by the vertices on the slider, and $E(x_{k},x_{k+1})$ corresponds to their mean velocities. The last two terms correspond to regularization on the contact forces and a penalty on the time spent in contact, where $T=Nh$. $k_{p},\,k_{v},\,k_{f},k_{T}\in\mathbb{R}_{>0}$ are cost parameters. We define the motion planning problem for a contact mode as minimizing (27) subject to (9), (16), (18), (19), (23), (26), and initial and final conditions on the state. This motion planning problem is non-convex due to the quadratic equality constraints arising from (9), (16) and (26). Specifically, the problem is a non-convex QCQP, which we relax into an SDP according to Subsection V-B.
For a non-contact mode $\mathcal{N}_{i},\,i=1,\ldots,N_{F}$, we minimize the following cost:
$\sum_{k=0}^{N-1}l_{p}\|p^{S}_{k+1}-p^{S}_{k}\|_{2}+l_{v}\|v^{S}_{k}\|_{2}^{2}+\sum_{k=1}^{N-1}l_{\phi}\frac{1}{\phi_{i}(x_{k})}$ (28)
where the first two terms model the arc length and velocity for the pusher trajectory respectively, and the last term encourages a positive distance between the slider and pusher when there is no contact. Since $\phi(x)>0$ when there is no contact, the last term can be represented as a Rotated Second-Order Cone (RSOC)constraint and is convex. $l_{p},\,l_{v},\,l_{\phi}\in\mathbb{R}_{>0}$ are cost parameters. We define the motion planning problem for a non-contact mode as minimizing the above cost function subject to the constraint (22a), the last row of (16), a constant slider pose, and initial and final conditions on the pusher position. Since these are all linear constraints, motion planning for a non-contact mode is naturally a convex program.
@@ VII-C Constructing the graph of convex sets
Finally, we formulate the motion planning problem as an SPP in a GCS. Let $G=(\mathcal{V},\mathcal{E})$ be the graph underlying a GCS as defined in Subsection V-A. For every contact mode $\mathcal{C}_{i},\,i=1,\ldots,N_{F}$ in the slider-pusher system, we add its corresponding SDP as a vertex $v$ to $\mathcal{V}$, in the sense that we let $\mathcal{X}_{v}$ and $l_{v}$ be the feasible set and cost function of the SDP, respectively. The point $x_{v}\in\mathcal{X}_{v}$ now corresponds to a trajectory of states and inputs for the slider-pusher system in mode $\mathcal{C}_{i}$.
Transitions between contact modes require the pusher to move within the non-contact modes, from one face of the slider to another. Since the pusher can visit every non-contact mode in each of these trajectories, we add to the graph a copy of all the non-contact modes for every possible contact-mode transition. Specifically, for any two contact modes $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ where $i,j\in\left\{1,\ldots,N_{F}\right\}$, $i\neq j$, we create a graph $G_{ij}=(\mathcal{V}_{ij},\mathcal{E}_{ij})$ where $\mathcal{V}_{ij}$ contains a vertex $v$ for each $\mathcal{N}_{k},\,k=1,\ldots,N_{F}$, where $\mathcal{X}_{v}$ and $l_{v}$ are the feasible set and cost function for the non-contact mode, and $\mathcal{E}_{ij}=\left\{(k,l)\mid\mathcal{Q}_{k}\cap\mathcal{Q}_{l}\neq\emptyset\right\}$ contains a bi-directional edge between all non-contact modes with intersecting regions. The vertices and edges of $\mathcal{G}_{ij}$ are added to $G$, as well as a bi-directional edge connecting the vertices corresponding to contact modes $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ to their respective non-contact modes to encode transition between contact and non-contact. An illustration is shown in Figure 2 (b). The total number of vertices in the graph is then $\mathcal{O}((N_{F})^{3})$, with $N_{F}$ contact modes and $N_{F}\binom{N_{F}}{2}$ non-contact modes.
We enforce continuity between the state trajectories on a path in the graph. Specifically, for each edge $e=(u,v)\in\mathcal{E}$ we enforce that the last state in the trajectory in vertex $u$ is equal to the first state in the trajectory in vertex $v$. Finally, we let the source vertex $s$ and target vertex $t$ be singleton sets that denote the desired starting and target states of the slider-pusher system, respectively. A feasible path $p$ through $G$ then has the interpretation as a continuous trajectory from the initial state to the target state, that consists of distinct trajectory segments for each contact mode represented by the vertices on the path, and the trajectory within each mode is determined by the points in the vertices.
@@ VII-D Additional remarks
First, we leverage (26) (which, for general rotation matrices looks like $R^{\intercal}R=I$) and add the dynamic constraints in (16) in both the world frame and the slider frame. While these constraints are obviously redundant for the QCQP due to (26), they yield different constraints in the semidefinite relaxation. Second, we use the initial and target rotations to generate tightening constraints. In (26) we parametrize the slider rotation by two variables in $\mathbb{R}^{2}$, where the start and target constraints on $\theta^{S}$ correspond to two points $r_{s}$ and $r_{t}$ on the unit circle. We define a hyperplane between these points in $\mathbb{R}^{2}$ with the normal vector pointing away from the origin, add the constraint that $r_{k}$ must be on the outside of this hyperplane. Again, this constraint is redundant for the QCQP where (26) holds, but is not reduntant in the semidefinite relaxation.
Additionally, we exploit structure in our problem that allows us to write a computationally cheaper semidefinite relaxation in (V-B). Specifically, the motion planning problem for a contact mode exhibits a band-sparse structure, in the sense that the non-convex quadratic constraints (9), (16), (26) only couple succeeding knot points $x_{k}$, $x_{k+1}$, and $u_{k}$, for $k=0,\ldots,N-1$. This allows us to form the relaxation (V-B) as an SDP with several smaller PSD matrices instead of a single, larger matrix. In principle, this does not include all the tightening constraints (4d) and yields a potentially weaker convex relaxation, but in practice, we find that the loss in tightness in negligible while significantly reducing the computational cost. For more details, the reader is referred to e.g.  [66].
@@ VIII Experiments
Figure 4: Our planner simultaneously reasons about both discrete mode switches and continuous motion. Here, an example of a planar pushing plan is shown for a rectangular slider geometry.
Slider SDP solve time Rounding time $\delta_{\text{round}}$
Box 5.89s (5.90s) 0.080s (0.065)s 5.95% (3.77%)
Tee 85.18s (86.56s) 0.11s (0.098s) s 9.74% (6.088%)
TABLE I: Statistics for planning planar pushing tasks with two different slider geometries, for 100 randomly chosen initial and final configurations. The box-shaped slider has 4 faces and the T-shaped slider has 8 faces. The reported values are mean values, with the median values shown in parenthesis.
@@
Figure 5: Our method is able to generate close-to globally optimal plans for pushing tasks with collision-free motion planning between contact modes. Here, three different pushing trajectories for a T-shaped slider are shown, stabilized with a feedback controller on a real robotic system.
@@
We show an example of a generated motion plan in Figure 4. The numerical experiments presented below can be reproduced using the code available at [67], and a video of the experiments can be seen at https://youtu.be/HBj7Vgoncxs. We choose the cost weights for contact modes to $k_{p}=10$, $k_{v}=100$, $k_{T}=1$, $k_{f}=1e5$, and $l_{p}=l_{v}=10$, $l_{\phi}=0.1$ for the non-contact modes, where the only tuning is choosing reasonable orders of magnitude for the cost parameters, so that the cost terms contribute approximately equally to the cost. For reference, both slider geometries have a max radius close to $0.25$ meters, and the plans are generated within a box with sides $0.6$ meters. We now evaluate the global optimality of the motion planner by generating 100 motion plans for two different slider geometries, with random initial and target configurations drawn from an uniform distribution, and upper-bound their respective optimality gaps according to (2). We use the conic solver Mosek [68] to solve the optimization problem on a laptop with an Apple M1 Max chip with 32GB of RAM. For the rectangular slider we achieve a success rate of $100\%$, that is, the rounding step is able to retrieve a feasible solution for all trajectories. For the T-shaped geometry, we achieve a success rate of $96\%$, where the remaining $4\%$ are due to numerical problems in the rounding. We show the average planning times, average rounding time, and average optimality gaps in Table I. We are able to achieve a mean upper bound on the optimality gap of less than $10\%$ and a median optimality gap of less than $6\%$ for both the slider geometries. We generate plans in a few seconds for the box-shaped slider , and in approximately $1.5$ minutes for the T-shaped slider. The mean solve time for the box-shaped slider is an order of magnitude lower than for the more T-shaped slider, which has twice as many faces, and hence generates a much bigger GCS problem.
@@
Next, we show statistics for the generated convex optimization problems for the two different slider geometries in Table II. It is clear that the size of the optimization problem grows quickly (although polynomially) with the number of faces on the slider. We believe that we can significantly reduce the size of the optimization problem, and thus reduce solve times, by better exploiting the heavily structured problem formulation, but leave this for future work.
@@
TABLE II: Statistics for the convex optimization problem that our methods formulates. The size of each PSD matrix is $15\times 15$, and $3$ knot points is used for each mode.

### v2

@@ VI-B The limit surface
and $\mu_{S}\in\mathbb{R}$ is the friction coefficient between the slider and table, $m_{S}\in\mathbb{R}$ is the slider mass, $g\in\mathbb{R}$ is the gravitional acceleration, $r\in\mathbb{R}$ is a characteristic distance, typically chosen as the max distance between a contact point and origin of frame $S$, and $c\in[0,1]$ is an integration constant that depends on the slider geometry, as detailed in [63] and [65]. From the principle of maximum dissipation, it can be shown that when the slider is sliding, the applied spatial contact force $F$ must lie on the limit surface, and the instantaneous spatial velocity $V$ must be perpendicular to the limit surface: $V\propto\nabla H(F)=DF$ , where ($\nabla$($\cdot$)∇ ( ⋅ ) denotes the gradient operator and $F$ and $V$ are measured in the same frame [48]. We note that this can be approximated as the linear constraint $V=DF$ and appropriately scaling $F$ as a post-processing step.
@@ VI-C Slider-pusher dynamics
$R=\begin{bmatrix}\prescript{W}{}{R}^{S}&0\\ 0&1\\ \end{bmatrix},\quad\prescript{W}{}{R}^{S}\in\mathrm{SO(2)}$
@@ VII-A Modes of the slider-pusher system
$r\coloneqq(c_{\theta},s_{\theta}),\quad\|r\|_{2}^{2}=1\iff\begin{bmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\\ \end{bmatrix}\in\mathrm{SO(2)}$ (16)
@@ VII-B Motion planning within a mode
We discretize the dynamics in (8) using forward Euler discretization. We represent a trajectory segment within each mode for the slider-pusher system by $N$ discrete knot points for the state and $N-1$ knot points for the input: $x_{0},x_{1},\ldots,x_{N}$ and $u_{0},u_{1},\ldots,u_{N-1}$. The dynamics are then $x_{k+1}=x_{k}+hg(x_{k},u_{k})$ for $k=0,\ldots,N-1$ with the chosen time step $h\in\mathbb{R}_{>0}$.
Our formulation allows any cost that is a combination (or subset) of arc length and trajectory energy for the slider and pusher, contact forces, and time spent in (and near) contact. The cost takes the following form:
$\sum_{k=0}^{N-1}$ $L(x_{k},x_{k+1})+E(x_{k},x_{k+1},u_{k})$
$+k_{f}h\|f_{k}\|_{2}^{2}+\sum_{k=0}^{N}\psi(x_{k})$ (17)
where the terms model the total trajectory arc length on $\mathrm{SE}(2)$ for the pusher and slider, the total kinetic energies for the pusher and slider trajectories, contact force regularization, and time spent near contact, respectively. Specifically
$L(x_{k},x_{k+1})$ $=k_{p^{P}}\|p^{P}_{k+1}-p^{P}_{k}\|_{2}$
$+k_{p^{S}}(1/N_{\nu})\sum_{i=1}^{N_{\nu}}\|p_{k+1}^{\nu_{i}}-p_{k}^{\nu_{i}}\|_{2}$ (18)
models the total arc lengths on $\mathrm{SE}(2)$ of the pusher and slider, and
$E(x_{k},x_{k+1},u_{k})$ $=k_{v^{P}}\|v^{P}_{k}\|_{2}^{2}$
$+k_{v^{S}}(1/N_{\nu})\sum_{i=1}^{N_{\nu}}\|v_{k}^{\nu_{i}}\|_{2}^{2}$ (19)
models the total kinetic energies of the trajectories for the pusher and slider. Recall that $p^{\nu_{i}}$ and $v^{\nu_{i}}$ refer to the position and velocity for slider vertex $i=1,\ldots,N_{\nu}$ in the world frame, where $N_{\nu}$ is the number of slider vertices. The total arc length of the slider trajectory on $\mathrm{SE}(2)$ is modelled through the mean arc length traversed by its vertices. Likewise, the energy of the slider trajectory is modelled through the mean squared velocity of its vertices. The constants $k_{p^{P}},\,k_{p^{S}},\,k_{v^{S}},\,k_{v^{P}}$ and $k_{f}\in\mathbb{R}_{>0}$ are cost weights for the arc lengths for the pusher and slider, trajectory energies for the pusher and slider, and contact force regularization.
The last term in (17) is a penalty on the distance between the pusher and the closest face of the slider. It is given by:
$\psi(x)=\max_{i=1,\ldots,N_{F}}\frac{hk_{T}}{1+\frac{1}{k_{\phi}}\phi_{i}(p^{P})}$ (20)
where $k_{T},\,k_{\phi}\in\mathbb{R}_{>0}$ are cost weights for the time in contact and for being “close” to the slider, respectively. This term is at its maximum when there is contact between the pusher and slider, and increases monotonically as the pusher gets closer to the slider. It has the property that $\psi(x)=hk_{T}$ when $\min_{i}\phi_{i}(x)=0$, i.e. the cost of contact between the pusher and slider is $k_{T}$ per second spent in contact. It also has the property that increasing $k_{\phi}$ increases the cost for being at a given distance to the object, i.e. $k_{\phi}$ controls the notion of “close”. Since the denominator is always positive, this function is a maximum over convex functions and is thus readily encoded through $N_{F}$ Rotated Second-Order Cone (RSOC)constraints [66].
@@
We define the motion planning problem for a contact mode as minimizing (17) subject to (5), (8), (10), (11), (15), (16), and initial and final conditions on the state. This motion planning problem is non-convex due to the quadratic equality constraints arising from (5), (8) and (16). Specifically, the problem is a non-convex QCQP, which we relax into an SDP according to Subsection V-B. Further, we define the motion planning problem for a non-contact mode as minimizing (17) subject to the constraint (14a), the last row of (8), a constant slider pose, and initial and final conditions on the pusher position. Since these are all linear constraints, motion planning for a non-contact mode is naturally a convex program.
@@ VII-C Constructing the graph of convex sets
Finally, we formulate the motion planning problem as an SPP in a GCS. Let $G=(\mathcal{V},\mathcal{E})$ be the graph underlying a GCS as defined in Subsection V-A. For every contact mode $\mathcal{C}_{i},\,i=1,\ldots,N_{F}$ in the slider-pusher system, we add its corresponding SDP as a vertex $v$ to $\mathcal{V}$, in the sense that we let $\mathcal{X}_{v}$ and $l_{v}$ be the feasible set and cost function of the SDP, respectively. The point $x_{v}\in\mathcal{X}_{v}$ now corresponds to a trajectory of length $N$ of states and inputs for the slider-pusher system in mode $\mathcal{C}_{i}$.
Transitions between contact modes require the pusher to move within the non-contact modes, from one face of the slider to another. Since the pusher can visit every non-contact mode in each of these trajectories, we add to the graph a copy of all the non-contact modes for every possible contact-mode transition. Specifically, for any two contact modes $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ where $i,j\in\left\{1,\ldots,N_{F}\right\}$, $i\neq j$, we create a graph $G_{ij}=(\mathcal{V}_{ij},\mathcal{E}_{ij})$ where $\mathcal{V}_{ij}$ contains a vertex $v$ for each non-contact mode $\mathcal{N}_{k},\,k=1,\ldots,N_{F}$ with $\mathcal{X}_{v}$ and $l_{v}$ taken as the feasible set and cost function for the mode, and $\mathcal{E}_{ij}$ contains a bi-directional edge between all non-contact modes with intersecting collision-free regions. The vertices and edges of $\mathcal{G}_{ij}$ are added to $G$, as well as a bi-directional edge connecting the vertices corresponding to contact modes $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ to their respective non-contact modes to encode transition between contact and non-contact. An illustration is shown in Figure 3 (b). The total number of vertices in the graph is then $\mathcal{O}((N_{F})^{3})$, with $N_{F}$ contact modes and $N_{F}\binom{N_{F}}{2}$ non-contact modes.
We enforce continuity between the state trajectories on a path in the graph. Specifically, for each edge $e=(u,v)\in\mathcal{E}$ we enforce that the last state in the trajectory in vertex $u$ is equal to the first state in the trajectory in vertex $v$. Finally, we let the source vertex $s$ and target vertex $t$ be singleton sets that denote the desired starting and target states of the slider-pusher system, respectively. A feasible path $p$ through $G$ then has the interpretation as a continuous trajectory from the initial state to the target state, that consists of distinct trajectory segments for each mode represented by the vertices on the path, and the trajectory within each mode is determined by the points in the vertices.
@@ VII-D Additional remarks
First, we leverage (16) (which, for general rotation matrices looks like $R^{\intercal}R=I$) and add the dynamic constraints in (8) in both the world frame and the slider frame. While these constraints are obviously redundant for the QCQP due to (16), they yield different constraints in the semidefinite relaxation. Second, we use the initial and target rotations to generate tightening constraints. In (16) we parametrize the slider rotation by two variables in $\mathbb{R}^{2}$, where the start and target constraints on $\theta^{S}$ correspond to two points $r_{s}$ and $r_{t}$ on the unit circle. We define a hyperplane between these points in $\mathbb{R}^{2}$ with the normal vector pointing away from the origin, and add the constraint that $r_{k}$ must be on the outside of this hyperplane for $k=1,\ldots,N$. Under the assumption that an optimal plan should always pick the shortest geodesic path between two elements of $\mathrm{SO(2)}$, this constraint is redundant for the QCQP where (16) holds, but is not redundant in the relaxation.
Additionally, we exploit structure in our problem that allows us to write a computationally cheaper semidefinite relaxation in (4). Specifically, the motion planning problem for a contact mode exhibits a band-sparse structure, in the sense that the nonconvex quadratic constraints (5), (8), (16) only couple succeeding knot points $x_{k}$, $x_{k+1}$, and $u_{k}$, for $k=0,\ldots,N-1$. This allows us to form the relaxation (4) as an SDP with several smaller PSD matrices instead of a single, larger matrix. In principle, this does not include all the tightening constraints (4d) and yields a potentially weaker convex relaxation, but in practice, we find that the loss in tightness is negligible while significantly reducing the computational cost. For more details, the reader is referred to e.g.  [69].
@@ VIII Experiments
Figure 4: Our planner simultaneously reasons about both discrete mode switches and continuous motion. Here, an example of a planar pushing plan with multiple mode switches is shown for a T-shaped slider geometry.
@@
Figure 5: Our method is able to generate close-to globally optimal plans for pushing tasks with collision-free motion planning between contact modes. Here, two different pushing trajectories for a T-shaped slider are shown, stabilized with a feedback controller on a real robotic system.
(a) Our method
(b) Contact-implicit method
Figure 6: A comparative example between our method and a contact-implicit method. Our method picks the shortest path around the object, while the baseline goes the longer way around twice, highlighting the fact that our method is capable of global reasoning. Our method also guarantees that the trajectory stays collision-free between contacts, while the baseline can be seen to clip the corners of the slider.
This section contains both numerical and hardware experiments. An open-source implementation of the proposed motion planning algorithm, as well as the code used to run the experiments, can be found at https://bernhardgraesdal.com/rss24-towards-tight-convex-relaxations/.
We show an example of a generated motion plan in Figure 4. We choose the cost weights for contact modes to $k_{p^{P}}=k_{p^{S}}=k_{v^{P}}=10$, $k_{v^{S}}=100$, $k_{f}=10$, $k_{T}=1$, and $k_{\phi}=0.1$ for the non-contact modes, where the only tuning that has been done is choosing reasonable orders of magnitude for the cost parameters, ensuring that the cost terms contribute approximately equally to the cost. We consider two different slider geometries, one simple box-shaped geometry and a T-shaped geometry, with the COM located according to a uniform mass distribution. For reference, both slider geometries have a maximum ”radius” close to $0.25$ meters, and the plans are generated within a box with sides $0.6$ meters.
VIII-A Planner performance
@@
We evaluate the global optimality of the motion planner by generating 100 motion plans for the two slider geometries, with random initial and target configurations drawn from an uniform distribution, and upper-bound their respective optimality gaps according to (2). We use the conic solver Mosek [70] to solve the optimization problem on a laptop with an Apple M1 Max chip with 32GB of RAM. For both slider geometries, we achieve a success rate of $100\%$, that is, the rounding step is able to retrieve a feasible solution for all the generated problem instances. We show the average planning times, average rounding time, and average optimality gaps in Table I. We achieve a mean upper bound on the optimality gap of around $10\%$, with a median optimality gap of less than $8\%$ for both slider geometries. We generate plans in a few seconds for the box-shaped slider, and in approximately $1.5$ minutes for the T-shaped slider. The mean solve time for the box-shaped slider is an order of magnitude lower than for the T-shaped slider, which has twice as many faces, and hence generates a much bigger GCS problem.
Slider SDP solve time Rounding time Optimality gap ($\delta_{\text{round}}$)
Box 7.05s (6.87s) 0.05s (0.05)s 8.33% (5.39%)
Tee 83.61s (80.12) 0.36s (0.014s) s 10.41% (7.47%)
TABLE I: Solve times, rounding times, and upper bounds on the global optimality gaps for planning planar pushing tasks with two different slider geometries, for 100 randomly chosen initial and final configurations. The box-shaped slider has 4 faces and the T-shaped slider has 8 faces. The reported values are mean values, with the median values shown in parenthesis.
@@
Next, we show statistics for the generated convex optimization problems for the two different slider geometries in Table II. It is clear that the size of the optimization problem grows quickly (although polynomially) with the number of faces on the slider. We believe that we can significantly reduce the size of the optimization problem and thus reduce solve times by better exploiting the heavily structured problem formulation, but leave this for future work.
@@
TABLE II: Statistics for the convex optimization problem that our methods formulates. The size of each PSD matrix is $15\times 15$, and $3$ knot points are used for each mode.
VIII-B Comparison with contact-implicit trajectory optimization
To compare our method with a state-of-the-art baseline for contact-rich planning, we select a direct, contact-implicit trajectory optimization method similar to those proposed in [32] and [34]. This method encodes contact modes implicitly using complementarity constraints that are relaxed and solves the problem as a sequence of NLPs with increasingly strict complementarity constraints until they are satisfied with equality. This method requires an initial guess, which we provide as a straight-line interpolation between the initial and target configuration. We generate another $100$ random initial and target configurations from an uniform distribution, and solve the planning problems with both methods.
Success Rate
Slider Our method Contact-implicit method
Box 100% 58%
Tee 100% 12%
TABLE III: A comparison of the success rate between our method and a contact-implicit trajectory optimization method. As our method is capable of global reasoning and does not rely on an initial guess, it has a much higher success rate compared to the baseline.
Table III presents a comparison between our method and the contact-implicit method. With a cost function that is specifically tuned to optimize the baseline’s performance, our method finds a solution in $100\%$ of the instances for both slider geometries. In contrast, the baseline often fails, finding a solution in $58\%$ of the instances for the box-shaped slider geometry and a mere $12\%$ for the T-shaped slider. For the T-shaped slider, the contact-implicit method appears to succeed only when the initial and target configurations require a small number of mode switches that are relatively close to the initial guess. This limitation is not surprising, as the baseline is a local method that relies heavily on its initial guess. The T-shaped slider has a significantly higher number of possible mode sequences and a more complex SDF compared to the box-shaped geometry, making it more difficult for the baseline to find a solution. On the other hand, our method solves all instances for both geometries. This highlights a key advantage of our approach: by reasoning on a global level, our method (empirically) always finds a solution, without relying on an initial guess.
The two transcriptions rely on different pre-specified parameters, such as the number of knot points and trajectory/mode durations. These differences make it challenging to conduct a rigorous comparison of the plan quality between the two methods. Instead, we provide a qualitative analysis to highlight some advantages of our method. We start by pointing out that the performance of the contact-implicit method in table III required significant tuning of the cost function and problem numerics. After this tuning however, the found plans empirically seem very reasonable. In contrast, our method requires minimal tuning and no initial guess while still finding high-quality plans for all given problem instances. Our method also provides a good metric for evaluating solution quality through the obtained upper bound on the optimality gap, unlike the baseline, where such an objective comparison is difficult.
Furthermore, the baseline requires careful consideration to avoid problems with unwanted contacts and penetration between discretization points. While these problems can certainly be addressed by e.g. finer trajectory discretization, velocity constraints, or tuning, these issues are naturally handled in our method, which guarantees that the entire path is collision-free between contacts, independently of the aforementioned factors. Additionally, while the baseline requires a fixed, pre-specified trajectory duration, our method can vary the trajectory length through the number of visited contact modes. However, it is important to note that our method still relies on a pre-specified trajectory duration in each contact mode. Figure 6 illustrates some of the mentioned qualitative differences through a comparison of trajectories generated by the two methods.
VIII-C Execution on real hardware